# MDA: An Interpretable Multi-Modal Fusion with Missing Modalities and Intrinsic Noise

### v2

@@ Abstract
Multi-modal fusion is crucial in medical data research, enabling a comprehensive understanding of diseases and improving diagnostic performance by combining diverse modalities. However, multi-modal fusion faces challenges, including capturing interactions between modalities, addressing missing modalities, handling erroneous modal information, and ensuring interpretability. Many existing researchers tend to design different solutions for these problems, often overlooking the commonalities among them. This paper introduces a novel multi-modal fusion framework that introduces Modal-Domain Attention (MDA), which constructs linear attention relationships between modalities through continuous attention, thereby achieving adaptive adjustment of the weights for each modality. It aims to facilitate the fusion of multi-modal information while allowing for the inclusion of missing modalities or intrinsic noise, thereby enhancing the representation of multi-modal data. We provide visualizations of accuracy changes and MDA weights by observing the modal fusion process, offering a comprehensive analysis of its interpretability. Experimental results indicate that the proposed MDA approach maintains state-of-the-art (SOTA) performance across various tasks on multiple public datasets. Worth noting is that we will make publicly available a multi-modal database for gastrointestinal diseases, which includes WLE, EUS, and report data. The visualization outcomes of MDA within the gastrointestinal disease dataset align closely with the conclusions of existing clinical studies regarding the dependency of different diseases on multiple modalities. The associated code will also be released to facilitate future research endeavors.
Introduction
@@
Recently, multi-modal fusion learning has demonstrated outstanding performance across various tasks (Huang et al. 2021), particularly in medical applications, as it integrates medical information from different modalities, offering comprehensive diagnostic evidence for healthcare professionals (Litjens et al. 2017). It has been employed in tasks such as disease-assisted diagnosis (Manh et al. 2022; Jiao et al. 2024) and prognostic forecasting (Ren et al. 2024a).
Despite efforts in medical multi-modal research, fundamental and challenging issues remain due to the complexity of multi-modal clinical data and the real-world application scenarios: Chanllenge1: The difficulty of multi-modal fusion arising from modality heterogeneity. Despite the individual strengths of each modality, integrating different modalities, including images and text, to improve disease diagnosis is still hindered by modality heterogeneity, which arises from the diverse data representations across different modalities. The cross-attention method is considered an efficient multi-modal fusion technique due to its ability to better capture the complex interrelationships between modalities, which have been applied to tasks such as image recognition diagnosis (Sun et al. 2023; Ma et al. 2023), object detection (Shen et al. 2024; Wen et al. 2024), segmentation (Ates, Mohan, and Celik 2023; Liu et al. 2024; Lin et al. 2023), and image fusion (Li and Wu 2024), among others. However, due to the adherence to the mutual attention between dual modalities, they are challenging to extend to tri-modal or even more modalities. Chanllenge2: The issue of modality missing in real-world scenarios. Missing modalities have garnered widespread interest in the medical fields (Wang et al. 2023; Ma et al. 2021). Methods for handling missing modalities can be broadly categorized into two types (Zhang et al. 2024): inference-based and non-inference-based. Inference-based methods (Dong and Sun 2023; Yang et al. 2022; Li et al. 2023; Ren et al. 2024b) strive to generate information for the missing modalities, however, they often overlook the assessment of the quality of the generated missing modalities and may lead to model collapse (Shumailov et al. 2024). For non-inference-based approaches (Liu et al. 2023; Xu et al. 2022b; Zhang et al. 2019; Wen et al. 2023) to incomplete multi-modal learning, they focus more on utilizing the available modalities to explore information and understand multi-modal data, but the information gap caused by missing modalities is difficult to bridge. Chanllenge3:Intrinsic noise. Multimodal learning that incorporates text may encounter the issue of intrinsic noise, which is more prevalent in the medical domain due to the heavy reliance on subjective annotations by clinical experts in diagnostic reports (Fusaroli et al. 2021). Intrinsic noise refers to the hidden inherent noise present in the text, which contradicts the knowledge conveyed by other modalities and may influence the final predictions. For example, it could occur when the report describes the echogenicity of a tumor in an EUS image as "hyperechoic" while the image itself reveals that the echogenicity is "heterogeneous." Selectively filtering accurate and advantageous multimodal information while avoiding errors that could compromise diagnostic accuracy remains an unresolved research challenge in multimodal feature fusion. Chanllenge4: The comprehensive interpretability of multi-modal fusion. Correctly interpreting AI model decisions has been a persistent challenge in achieving interpretability in multi-modal fusion (Huff, Weisman, and Jeraj 2021; Salahuddin et al. 2022). Currently, the interpretability of medical-assisted diagnostic models primarily focuses on generating heatmaps (Selvaraju et al. 2017; Wang et al. 2020) to visualize the network’s attention and enhance model interpretability (Manh et al. 2022; Wu et al. 2023). However, heatmaps are visualizations of network parameters and cannot explain modality specificity in different cases, i.e., they cannot reflect modality weights for different cases. In the context of diagnosing gastrointestinal diseases, different cases may focus on different modalities (Jacobson et al. 2023), and the inability to capture the inconsistency between modalities or decisions that contradict clinical experience may lack credibility. The interpretability of fine-grained features is undoubtedly crucial. However, the interpretability of modalities, that is, discussing the gold standard modalities for diagnosing various diseases, is equally important  (Richardson et al. 1995; Chandler et al. 2019; Jacobson et al. 2023). To address the aforementioned challenges, we propose a novel modality fusion approach wherein we introduce MDA. The objective of MDA is to compute the saliency of a particular modality while considering the interactive influences of all other modalities, thereby achieving a comprehensive understanding of the multimodal information within the overall scene. This is primarily achieved through the implementation of a novel continuous attention algorithm, which constructs linear attention relationships between multi-modal features, thereby dynamically adjusting the attention allocation between different modalities for the thorough integration of multi-source heterogeneous information (as shown in Fig. 1). Furthermore, this approach ensures the scalability of the fusion framework, offering a new solution for scalable multimodal fusion. Additionally, it pioneers the exploration of the macroscopic interpretability of multimodal learning.
@@
Figure 1: A unified multi-modal learning strategy involves learning with different multi-modal configurations. (a) Train and test with full modality. (b) The model will reduce its attention when learning with missing modalities or intrinsic noise.
  • We propose a novel multi-modal fusion framework that integrates MDA, which constructs linear relationships between multi-modal information through continuous attention, thereby achieving adaptive control over the weights of multiple modalities and demonstrating modality scalability.
  • MDA provides a unified solution to several challenges in multi-modal fusion, including multi-modal integration, modality missing, and learning with intrinsic noise.
  • Supplementary experiments indicate that we achieve SOTA performance across multiple tasks on various public datasets. Additionally, we provide a publicly accessible multi-modal gastrointestinal stromal tumor dataset.
  • Based on MDA, we comprehensively analyze the sources of multi-modal efficacy from a macroscopic perspective (across different diseases) and a microscopic perspective (within individual cases) while providing interpretability to the model from a clinical experience perspective.
@@
Figure 2: Overview of the proposed network framework. The uni-modal training involves building pre-trained models for multi-disease classification across different modalities, with a dedicated classifier assigned to each disease. The work introduces the weight calculation for each scenario under the proposed multi-modal attention module.
Methodology
This study aims to enhance collaborative knowledge learning capabilities in various modalities, including different image modalities, text modality, audio modality, and others. Considering there are ${M}$ image modalities ${I}$, ${P}$ text modalities ${T}$, and ${Q}$ audio modalities ${A}$, the total number of modalities is ${N(N=M+P+Q)}$. The multi-modal data ${\mathcal{X}}$ can be represented as $\mathcal{X}=\{\{\mathbf{I}_{j}\}_{j=1}^{M},\{\mathbf{T}_{k}\}_{k=1}^{P},\{\mathbf{A}_{l}\}_{l=1}^{Q}\}$ and the predicted labels as ${Y}$.
An overview of the proposed method is shown in Fig. 2. we first construct individual classification tasks ${\mathcal{T}_{j}=\{\mathbf{I}_{j},Y\}}$ for modal ${j}$, the single-task definitions for other modalities are similar. Then, we feed the extracted features into the MDA to calculate the adaptive weights for each modality. Finally, we perform multi-modal feature fusion based on the inter-modality weights and feed the fused features into the classifier to accomplish the classification task. The entire multi-modal classification task be defined as ${\mathcal{T}=\{\mathcal{X},Y\}}$, ${Y}$ is represented as a one-hot vector for this multi-class task. Next, we will define the processing approaches for different modalities.
Image modalities
@@
Pre-trained convolutional neural networks (CNNs) is utilized as feature extractors for image modalities, such as ResNet. Take image modality $\mathbf{I}_{j}$ as an example, the vectorized latent features ${\mathbf{f}_{j}}$ obtained from the feature extractors ${\mathbf{G}_{j}}$ are denoted as:
@@
Text modality
We employ BERT (Devlin et al. 2018) to encode the input text. BERT learns contextual representations of words or subwords in a text by using a self-attention mechanism. This is consistent with the self-attention module in image feature extraction. To adapt BERT to our specific task, we unfroze the last four layers of BERT for training. The self-attention feature extraction $\mathbf{\Phi}_{k}$ representation for textual reports $\mathbf{T}_{k}$ is as follows:
@@
Audio modality
We utilize Mel-frequency cepstral coefficients (MFCCs) (Tzanetakis and Cook 2002) as the representation for the audio modality. MFCCs have gained extensive application in the field of audio and speech processing due to their ability to mimic human auditory perception, facilitate dimensionality reduction of features, exhibit robustness to variations, and demonstrate broad applicability. We also applied self-attention to the extracted audio features, take audio modality ${\mathbf{A}_{l}}$ as an example, and its feature representation $\mathbf{\Phi}_{l}$ obtained after processing through the self-attention mechanism is denoted as:
@@
Continuous attention mechanism
Existing multi-modal attention mechanisms are typically designed for bi-modal fusion, focusing on point-to-point interactions between modalities. However, extending these methods to handle three or more modalities leads to computational complexity and a lack of consideration for interactions between modalities. In contrast, MDA aims to compute the attention of each modality while considering the influences of all other modalities, enabling a comprehensive understanding of multi-modal information. It achieves this by employing continuous attention mechanisms to dynamically adjust the attention across all modalities and to establish linear relationships between attentions, facilitating rapid and dynamic adjustments. The continuous attention mechanism employed in MDA constructs linear attention relationships among multiple modalities. The specific computational approach is illustrated as Eq. 5, where $\mathbf{\Theta}_{i}^{N-1}$ represents the attention value computed between modality $i$ and the other $N-1$ modalities, $R$ denotes the ReLU activation.
$\mathbf{\Theta}^{e}_{i}=\begin{cases}\operatorname{R}\left(\frac{Q_{i}\cdot(K_{e})^{T}}{\sqrt{d}}\right)\cdot V_{i}^{e},e=1\\ \operatorname{R}\left(\frac{\mathbf{\Theta}^{e-1}_{i}\cdot(K_{e})^{T}}{\sqrt{d}}\right)\cdot V_{i}^{e},1\leq e\leq N-1\end{cases}$ (5)
MDA establishes a linear relationship between multi-modal attentions through continuous attention calculations, allowing for infinite expansion without being limited to handling only two or three modalities. This scalability is crucial for applications that require continuous integration of new modalities over time, particularly in medical imaging, where the collection of more modal data is necessary to provide accurate and trustworthy judgments.
Objective function
For the classification training of single-modality model $i$, given the features ${\mathbf{\Phi}_{i}}$ obtained after the self-attention module, we employ a multi-layer perceptron classifier ${\mathcal{C}_{s}}$ for disease prediction. The model training is guided by the cross-entropy loss, defined as Eq. 6. For the multi-modal classification training, the classification loss is defined as Eq. 7, where ${\mathcal{C}_{f}}$ stands for multi-modal classifier.
@@
Training setting
We have directly developed a model that operates under conditions of missing modalities and intrinsic noise. We processed the individual modality features obtained to simulate these scenarios. Specifically, within each batch of data, we randomly selected ${(N+P)}$ samples and subjected them to the following treatments: $N$ modality’s features were sequentially set to 1 to simulate missing modalities, and one random attribute in $P$ textual modality was replaced with another random label to simulate intrinsic noise.
@@
Datasets
We aim to develop a general multi-modal fusion model that can effectively handle modality heterogeneity, adapt to varying degrees of missing modalities, filter out inconsistent and intrinsic noise, and mitigate their negative impact on the results. To this end, we conducted experiments using Four different multi-modal datasets, encompassing diverse dimensions such as medical data, natural images, movies, and audio, to assess the generality of the proposed model comprehensively.
Gastrointestinal disease dataset We compiled a multi-center gastrointestinal disease dataset comprising three modalities: Endoscopic Ultrasound (EUS), White Light Endoscopy (WLE), and imaging reports. The dataset was gathered from two separate centers (***hospital and ***hospital) and captured using Olympus UM-2R and UM-3R equipment, totaling 421 cases spanning eight distinct categories (Gastrointestinal Stromal Tumor (GIST), Gastric Leiomyoma (GLM), Neuroendocrine Tumor (NET), Ectopic Pancrea (EP), Lipoma, Gastrointestinal Schwannomas (GS), Pneumatosis Cystoid (PC), and Inflammatory Fibroid Polyp (IFP)). Among these patients, the gender distribution is 53% male to 47% female, with an age range spanning from 27 to 78 years. The imaging reports offer detailed descriptions of lesion characteristics observed within the two imaging modalities. The detailed data quantities for each category are shown in Table 1.

### v3

# MDA: An Interpretable and Scalable Multi-Modal Fusion under Missing Modalities and Intrinsic Noise Conditions

@@ Abstract
Multi-modal learning has shown exceptional performance in various tasks, especially in medical applications, where it integrates diverse medical information for comprehensive diagnostic evidence. However, there still are several challenges in multi-modal learning, 1. Heterogeneity between modalities, 2. uncertainty in missing modalities, 3. influence of intrinsic noise, and 4. interpretability for fusion result. This paper introduces the Modal-Domain Attention (MDA) model to address the above challenges. MDA constructs linear relationships between modalities through continuous attention, due to its ability to adaptively allocate dynamic attention to different modalities, MDA can reduce attention to low-correlation data, missing modalities, or modalities with inherent noise, thereby maintaining SOTA performance across various tasks on multiple public datasets. Furthermore, our observations on the contribution of different modalities indicate that MDA aligns with established clinical diagnostic imaging gold standards and holds promise as a reference for pathologies where these standards are not yet clearly defined. The code and dataset will be available.
1 Introduction
@@
Recently, multi-modal learning has demonstrated outstanding performance across various tasks [13], particularly in medical applications, as it integrates information from different modalities, offering comprehensive diagnostic evidence for healthcare professionals [26]. It has been employed in tasks such as disease-assisted diagnosis [30, 16] and prognostic forecasting [34].
@@
Figure 1: A unified multi-modal learning strategy involves learning with different multi-modal configurations. (a) Train and test with full modality. Different modalities receive equal attention (b) The model will reduce its modal-domain attention when learning with missing modalities or intrinsic noise.
Despite efforts in multi-modal medical research, fundamental and challenging issues remain due to the complexity of multi-modal clinical data and real-world application scenarios: Chanllenge1: Multi-modal heterogeneity challenges fusion. Modality heterogeneity includes information redundancy and complementary misalignment, which pose challenges for fusion. The greater the number of modalities, the more difficult the fusion process becomes. Various fusion methods have been proposed, initially based on multi-scale transformations  [17, 31] and sparse representations  [25, 42], and later using deep learning techniques [49, 50, 52]. However, these methods fail to effectively exploit modality synergy, often relying on simple concatenation of latent features, allocating the same attention to each modality, which limits the potential of multi-modal integration. Recently, multi-modal attention mechanisms have been introduced [38, 29], but they primarily focus on mutual attention between two modalities. A key limitation of the above methods is their inability to dynamically allocate attention based on the importance of each modality to the final diagnosis, especially in the case of multiple heterogeneous multi-modal data, as shown in Fig. 1 (a).
Figure 2: Existing methods vs. MDA for missing modality. Generation-based methods  [6] add specific models (Gen-1, Gen-2) for missing modalities, similar to prompt-based methods  [22] that add different prompts (Pr-1,Pr-2, Pr-3), complexifying network parameters intensify training difficulty. This work builds linear attention relationships between modalities, adaptively adjusting their weights in real-time, and more efficiently handling various scenarios involving missing modalities.
@@
Figure 3: Overview of the proposed network framework. The uni-modal training involves building pre-trained models for multi-disease classification across different modalities, with a dedicated classifier assigned to each disease. The work introduces the weight calculation for each scenario under the proposed modal-domain attention module.
Chanllenge2: The issue of modality missing in real-world scenarios. Missing modality is an inevitable issue in medical multi-modal learning. Several generate-based methods [48, 24, 35] are dedicated to designing distinct generative models for each missing modality to achieve full modality effects, as shown in Fig. 2 (a). Recently, a series of approaches including MAPs [22, 7, 15] fine-tuning the Visual-Language model by constructing specialized prompts for each missing modality to address the issue, as shown in Fig. 2 (b). While these prompt-based methods offer a more flexible and scalable solution than generate-based methods. However, the inherent complexity of designing and maintaining a large number of generative networks or prompts increases with the number of modalities [12]. Moreover, the assumption that a dedicated generative network can effectively handle each missing modality or prompt may not always hold, especially in scenarios where the relationships between modalities are complex and non-linear.
Chanllenge3: Learning with intrinsic noise. As the field of multi-modal research gains momentum, the intra-modal cross-verification of information presents innovative perspectives for noise learning [27], rather than focusing on label noise [18, 45, 32]. The exploration of intrinsic noise within text data remains limited in multi-modal research. Intrinsic noise, that is, discrepancies between textual information and clinical realities, can introduce contradictions during modality fusion. This phenomenon is particularly pronounced in the medical domain, where diagnostic reports are often laden with subjective annotations made by clinical experts, as highlighted by Fusaroli [9]. Exploiting the complementary differences between modalities to mitigate the impact of intrinsic noise may serve as an effective solution in multi-modal fusion.
Chanllenge4: The comprehensive interpretability of multi-modal fusion. Currently, techniques such as heat maps are widely used to improve the interpretability of models, allowing clinicians to gain direct insights into the regions underlying diagnostic decisions [30, 46]. However, these methods are limited to providing explanations within the confines of individual modalities and do not extend to cross-modal scenarios. The interactive interpretability between multi-modal lacks a unified paradigm, which is precisely what is needed in clinical practice. For instance, comprehensive diagnosis of gastrointestinal tumors in clinical settings often utilizes multi-modal imaging techniques such as EUS, WLE, CT, and MRI, with each modality being sensitive to different disease types  [14]. Some diseases have established diagnostic gold standards, such as WLE for lipomas. However, many other conditions still lack definitive diagnostic criteria. The medical community is increasingly focused on establishing diagnostic gold standards, aiming to standardize procedures and reduce the number of imaging modalities required  [36, 5, 14]. Therefore, exploring how to construct a paradigm for interactive interpretability between multi-modalities has become an urgent task in the field of medical image analysis.
We argue that the essence of the challenges above is uniform, focusing on how to efficiently and flexibly process and integrate multi-modal data, each carrying distinct information. We aim to address the above challenges by adjusting the weight of each modality in the final decision-making process. Guided by this, we introduce the Modal-Domain Attention (MDA), which employs an innovative continuous attention algorithm, establishing a linear attention relationship between multi-modal features, flexibly allocating attention to different modalities: it reduces the attention to modalities with lower relevance to the final decision, to missing modalities, and to those containing intrinsic noise, as illustrated in Fig. 1 (b), and can more efficiently handle various scenarios involving missing modalities, as illustrated in Fig. 2 (c). Our contributions are as follows:
  • MDA can adaptively adjust the weight of each modality in the final decision-making process and has achieved SOTA performance across multiple multi-modal datasets, demonstrating that this strategy can effectively enhance performance in multi-modal tasks.
  • MDA can adjust the attention for each modality in decision-making, reducing the emphasis on low-relevance information and minimizing the impact of missing modalities and intrinsic noise on outcomes.
  • Based on MDA, this study comprehensively analyzes the sources of multi-modal efficacy from both macroscopic (across different diseases) and microscopic perspectives (within individual cases), while pioneering in exploring the interpretability of multi-modal learning as the gold standard for clinical diagnostics, thereby promoting the precision and standardization of medical diagnostics for standards-lacking diseases.
2 Methodology
This study introduces MDA, an adaptable multi-modal fusion strategy. MDA fortifies against the missing of any modality and reduces intrinsic noise influence by adjusting the modal contributions to decision-making. These contributions index the decision’s reliance on each modality. The overview of the method is shown in Fig. 3. Given ${M}$ image modalities ${I}$, ${P}$ text modalities ${T}$, ${Q}$ audio modalities ${A}$. The total number of modalities is ${N(N=M+P+Q)}$. The multi-modal dataset ${\mathcal{X}}$ can be represented as $\mathcal{X}=\{\{\mathbf{I}_{j}\}_{j=1}^{M},\{\mathbf{T}_{k}\}_{k=1}^{P},\{\mathbf{A}_{l}\}_{l=1}^{Q}\}$, with labels ${Y}$. Then Individual classification tasks ${\mathcal{T}_{j}=\{\mathbf{I}_{j},Y\}}$ are defined for each modal, with analogous definitions for others. Finally, features are processed by MDA to determine adaptive weights. The final classification is achieved by fusing modal features based on these weights and inputting them into a classifier. The multi-modal classification task is defined as ${\mathcal{T}=\{\mathcal{X},Y\}}$, where ${Y}$ is a one-hot vector for multi-class classification. Subsequent sections detail the processing of each modality.
2.1 Image modalities
@@
Pre-trained convolutional neural networks (CNNs) are utilized as feature extractors for image modalities. Take image modality $\mathbf{I}_{j}$ as an example, the vectorized latent features ${\mathbf{f}_{j}}$ obtained from the feature extractors ${\mathbf{G}_{j}}$ are denoted as:
@@
2.2 Text modalities
We employ BERT [8] to encode the input text. BERT learns contextual representations of words or subwords in a text by using a self-attention mechanism. This is consistent with the self-attention module in image feature extraction. To adapt BERT to our specific task, we unfroze the last four layers of BERT for training. The self-attention feature extraction $\mathbf{\Phi}_{k}$ representation for textual reports $\mathbf{T}_{k}$ is as follows:
@@ 2.2 Text modalities
2.3 Audio modalities
We utilize Mel-frequency cepstral coefficients (MFCCs) [40] as the representation for the audio modality. MFCCs have gained extensive application in the field of audio and speech processing due to their ability to mimic human auditory perception, facilitate dimensionality reduction of features, exhibit robustness to variations, and demonstrate broad applicability. We also applied self-attention to the extracted audio features, taking audio modality ${\mathbf{A}_{l}}$ as an example, and its feature representation $\mathbf{\Phi}_{l}$ obtained after processing through the self-attention mechanism is denoted as:
@@ 2.3 Audio modalities
2.4 Continuous attention mechanism
Existing multi-modal attention mechanisms are typically designed for bi-modal fusion, focusing on point-to-point interactions between modalities, overlooking the collective influence of all other modalities. We assert that in real-world multi-modal data processing, the information or semantics of one modality are inherently intertwined with those of all other modalities, necessitating a more holistic and scalable fusion technique. In contrast, MDA aims to compute the attention for each modality while taking into account the influence of all other modalities, thereby achieving a comprehensive understanding of multi-modal information. This approach also enhances the scalability of the expansion of modalities in any given scenario. It achieves this by employing a continuous attention mechanism that constructs linear attention relationships among multi-modalities, allowing it to dynamically adjust the contribution of each modality to the decision-making process. This reduces the impact of modalities containing invalid or erroneous information, thereby enhancing the overall accuracy. The specific computational approach is illustrated as Eq. 5, where $\mathbf{\Theta}_{i}^{N-1}$ represents the attention map computed between modality $i$ and the other $N-1$ modalities, which is also known as the MDA map $\mathbf{\Theta}_{i}$. $R$ denotes the ReLU activation.
$\mathbf{\Theta}^{e}_{i}=\begin{cases}\operatorname{R}\left(\frac{Q_{i}\cdot(K_{e})^{T}}{\sqrt{d}}\right)\cdot V_{i}^{e},e=1\\ \operatorname{R}\left(\frac{\mathbf{\Theta}^{e-1}_{i}\cdot(K_{e})^{T}}{\sqrt{d}}\right)\cdot V_{i}^{e},1\textless e\leq N-1\end{cases}$ (5)
MDA allows for infinite expansion without being limited to handling only two or three modalities, and the network complexity remains consistently at a low level. This scalability is crucial for applications that require continuous integration of new modalities over time, particularly in medical imaging, where the collection of more modal data is necessary to provide accurate and trustworthy judgments.
2.5 Objective function
For the classification training of single-modality model $i$, given the features ${\mathbf{\Phi}_{i}}$ obtained after the self-attention module, we employ a multi-layer perceptron classifier ${\mathcal{C}_{s}}$ for disease prediction. The model training is guided by the cross-entropy loss, defined as Eq. 6. For the multi-modal classification training, the classification loss is defined as Eq. 7, where ${\mathcal{C}_{f}}$ stands for the multi-modal classifier.
@@ 2.5 Objective function
2.6 Training setting
We have directly developed a model that operates under conditions of missing modalities and intrinsic noise. We processed the individual modality features obtained to simulate these scenarios. Specifically, we randomly select a certain proportion of samples from each batch of data and apply the following treatments: 1. randomly set the features of one modality to 1 to simulate the arbitrary missing modality, and 2. replace an attribute within the text modality with another random description to simulate intrinsic noise. 3. The rates of missing modality or inherent noise in the training and test sets is identical.
@@ 2.6 Training setting
3 Datasets
We conducted experiments using four different multi-modal datasets, encompassing diverse dimensions such as medical data, natural images, movies, and audio, to assess the generality of the proposed model comprehensively.